# Extremal distributions of partially hyperbolic systems: the Lipschitz threshold

### 01 Apr 2026 v1

@@ 4.1. Adapted charts and templates
    let $F_{x}:=(\imath_{f(x)})^{-1}\circ f\circ\imath_{x}=(F_{x,1},F_{x,2},F_{x,3})$; then, there exist quadratic polynomials $P_{x}^{s}(z)=\alpha_{x}z+\beta_{x}^{s}z^{2}$, $P_{x}^{u}(z)=\alpha_{x}z+\beta_{x}^{u}z^{2}$, such that for $\xi\in(-1,1)$ and $\eta\in(-(\lambda_{x}^{u})^{-1},(\lambda_{x}^{u})^{-1})$, we have
@@
    (4) $\alpha_{x}:=\partial_{13}F_{x,2}(0,0,0),\quad\forall\,x\in M.$
@@
To obtain the desired polynomials as stated in Proposition 4.2, two additional cohomological equations must be solved. We focus on the case of $r_{x}^{u}$; the other case follows similarly. We decompose $r_{x}^{u}(\xi)$ as $r_{x}^{u}(\xi)=P_{x}^{u}(\xi)+\hat{r}_{x}^{u}(\xi)$, where $P_{x}^{u}(\xi)=\alpha_{x}\xi+\beta_{x}^{u}\xi^{2}$ and $\hat{r}_{x}^{u}(\xi)=O(\xi^{3})$. As in [27, Claim 3.5], the required chart adjustment involves solving the following cohomological equation for $u$:
@@ 4.2. The Foulon–Hasselblatt cocycle
We define the (twisted) Foulon–Hasselblatt cocycle $\alpha$ to be the map
$M\times\mathbb{Z}\ni(x,n)\mapsto\alpha_{x}(n):=\left\{\begin{array}[]{ll}\sum_{\ell=0}^{n-1}\lambda_{x}(\ell)\alpha_{f^{\ell}(x)},&n\geq 0,\\ -\sum_{\ell=1}^{-n}\lambda_{x}(-\ell)\alpha_{f^{-\ell}(x)},&n<0,\end{array}\right$.
with $\alpha_{x}$ as in (4). Note that the twist is given by the cocycle $\lambda$ defined above.
@@
The Foulon–Hasselblatt cocycle $\alpha$ is a twisted coboundary if
$\alpha_{x}=\lambda_{x}\beta(f(x))-\beta(x),\quad\forall\,x\in M,$
@@
The twisted cohomology class of the Foulon–Hasselblatt cocycle is independent of the choice of adapted charts. In other words, if $\{\tilde{\imath}_{x}\}_{x\in M}$ is another family of adapted charts, and $\tilde{\alpha}$ denotes the associated Foulon–Hasselblatt cocycle, then the difference $\tilde{\alpha}-\alpha$ is a twisted coboundary, i.e.,
$\tilde{\alpha}_{x}-\alpha_{x}=\lambda_{x}\beta(f(x))-\beta(x),\quad\forall\,x\in M,$
@@
$\begin{bmatrix}\partial_{1}\tilde{F}_{x,1}&\partial_{2}\tilde{F}_{x,1}\\ \partial_{1}\tilde{F}_{x,2}&\partial_{2}\tilde{F}_{x,2}\end{bmatrix}(0,0,\eta)$ $=\begin{bmatrix}\lambda_{x}^{s}&0\\ \tilde{\alpha}_{x}\eta+\tilde{\beta}_{x}^{s}\eta^{2}&\lambda_{x}^{c}\end{bmatrix}$.
@@
$=(\tilde{\alpha}_{x}+\lambda_{x}^{c}\tau_{x})\xi\eta+H.O.T.$
@@
$\Phi_{f(x)}\circ\tilde{F}_{x}(\xi,0,\eta)$ $=\Phi_{f(x)}(\lambda_{x}^{s}\xi+H.O.T.,\alpha_{x}\xi\eta+H.O.T.,\lambda_{x}^{u}\eta+H.O.T.)$
$=(*,(\alpha_{x}+\tau_{f(x)}\lambda_{x}^{s}\lambda_{x}^{u})\xi\eta+H.O.T.,*).$
@@
(7) $\tilde{\alpha}_{x}-\alpha_{x}=\lambda_{x}^{s}\lambda_{x}^{u}\tau_{f(x)}-\lambda_{x}^{c}\tau_{x}=\lambda_{x}\beta(f(x))-\beta(x),$
@@
$\alpha_{x}=\lambda_{x}\beta(f(x))-\beta(x),\quad\forall\,x\in M.$
We look for a family $\{\tilde{\imath}_{x}\}_{x\in M}$ such that the associated family of (quadratic) polynomials $\tilde{P}_{x}^{*}$ have a vanishing linear coefficient, i.e., $\tilde{\alpha}_{x}=0$. As above, we let $\Phi_{x}:=\tilde{\imath}_{x}^{-1}\circ\imath_{x}\colon(\xi,t,\eta)\mapsto(\tilde{\xi}_{x},\tilde{t}_{x},\tilde{\eta}_{x})$ be the change of coordinates; by (7), we then have
$\tilde{\alpha}_{x}=\alpha_{x}+\lambda_{x}\lambda_{f(x)}^{c}\partial_{13}\tilde{t}_{f(x)}(0,0,0)-\lambda_{x}^{c}\partial_{13}\tilde{t}_{x}(0,0,0).$
To ensure that $\tilde{\alpha}_{x}=0$, it suffices to solve $\lambda_{x}^{c}\partial_{13}\tilde{t}_{x}(0,0,0)=\beta(x)$; we can then take
@@
The non-triviality of the Foulon–Hasselblatt cocycle $\alpha$ is an obstruction to the regularity of the distribution $E^{s}\oplus E^{u}$.
@@
If $E^{s}\oplus E^{u}$ is $C^{1}$ along the leaves of the unstable foliation $\mathcal{W}^{u}$, then $\alpha$ is a twisted coboundary. The same conclusion holds if $E^{s}\oplus E^{u}$ is $C^{1}$ along the leaves of the stable foliation $\mathcal{W}^{s}$.
@@
$\alpha_{x}$ $=\lambda_{x}^{s}\lambda_{x}^{u}(\mathcal{T}_{f(x)}^{s})^{\prime}(0)-\lambda_{x}^{c}(\mathcal{T}_{x}^{s})^{\prime}(0)$
@@ 4.3. Bootstrap of regularity
Note that $f$ is dynamically coherent and center bunched, hence the distribution $E^{s}\oplus E^{u}$ is uniformly $C^{1}$ along $\mathcal{W}^{c}$. By Journé’s lemma [42], it suffices to show that $E^{s}\oplus E^{u}$ is uniformly $C^{1}$ along local manifolds $\mathcal{W}^{u}_{\mathrm{loc}}(x)$ and $\mathcal{W}^{s}_{\mathrm{loc}}(x)$, $x\in M$, respectively. By construction, this is equivalent to showing that the templates $\mathcal{T}^{s}_{x}$ and $\mathcal{T}^{u}_{x}$ are uniformly $C^{1}$. We prove it for $\mathcal{T}^{s}_{x}$ here, the proof for $\mathcal{T}^{u}_{x}$ is similar.
@@
$\alpha_{x}=\lambda_{x}^{s}\lambda_{x}^{u}(\mathcal{T}_{f(x)}^{s})^{\prime}(0)-\lambda_{x}^{c}(\mathcal{T}_{x}^{s})^{\prime}(0)=\lambda_{x}\beta(f(x))-\beta(x)$
@@
$\alpha^{\prime}_{x}:=\gamma(x)\alpha_{x}=\gamma(x)(\lambda_{x}\beta(f(x))-\beta(x))=\beta(f(x))\gamma(f(x))-\beta(x)\gamma(x)$
@@
By Lemma 4.11, the distribution $E^{s}\oplus E^{u}$ is $C^{1}$. This implies that $x\mapsto\alpha_{x}$ is a twisted coboundary by Lemma 4.10. Then by Lemma 4.9, up to replacing the charts $\{\imath_{x}\}_{x\in M}$ by new charts $\{\tilde{\imath}_{x}\}_{x\in M}$, we can assume that the Foulon–Hasselblatt cocycle vanishes identically. In other words, for all $x\in M$, the degree-one coefficient $\alpha_{x}=(P_{x}^{s})^{\prime}(0)=0$ of the polynomial $P_{x}^{s}$ vanishes; consequently, $P_{x}^{s}(\eta)=\beta_{x}^{s}\eta^{2}$. By (5), for $x\in M$ and $\eta\in(-1,1)$, we have
@@ 5.3. Smoothness of $E^{s}\oplus E^{u}$
We then show that if $E^{s}\oplus E^{u}$ is Lipschitz then it is actually $C^{\infty}$. Recall that we have fixed a non-trivial $f$-invariant open accessibility class $U\subsetneq M$ of $f$. By applying Proposition 4.12, in order to show that $E^{s}\oplus E^{u}$ is $C^{\infty}$ on $U$, it suffices to check that
@@
The map $f_{\bar{U}}$ is $\infty$-bunched.
@@
Let $\nu$ be any $f$-invariant ergodic measure supported on $\bar{U}$, and let $x\in\bar{U}$ be a $\nu$-typical point. If $x\in U$, then by Corollary 3.3, the center Lyapunov exponent satisfies
@@ 6.1. Proof of Theorem A
For (3), Proposition 5.1 implies that either $f$ is accessible or $E^{s}\oplus E^{u}$ is integrable. Note that $\lambda^{c}_{x}$ is a coboundary if and only if $f$ is center isometric, i.e., there is a continuous metric such that $\lambda^{c}_{x}=1$, for all $x\in M$. In the accessible case, $\lambda^{c}_{x}$ is a coboundary by Corollary 3.3. If $E^{s}\oplus E^{u}$ is integrable, then by Proposition 2.7 the map $f$ is either center isometric or topologically conjugate to an Anosov automorphism of ${\mathbb{T}}^{3}$. In the latter case $f$ is Anosov, and $\lambda^{c}$ is cohomologous to a constant by [30, Theorem 1.1].
For (1), center isometric partially hyperbolic systems are dynamically coherent by [55, Theorem 7.5], while Anosov diffeomorphisms on ${\mathbb{T}}^{3}$ are always dynamically coherent [53].
For (2), if $f$ is center isometric and accessible, then $E^{s}\oplus E^{u}$ is $C^{\infty}$ by Proposition 4.12. If $E^{s}\oplus E^{u}$ integrates to a foliation $\mathcal{W}^{su}$, then the leaves of $\mathcal{W}^{su}$ are $C^{\infty}$ by Lemma 2.10. Since $f$ is $\infty$-bunched, the holonomy of $\mathcal{W}^{su}$ along $\mathcal{W}^{c}$ is uniformly $C^{\infty}$. Journé’s lemma [42] then yields that $\mathcal{W}^{su}$ is a $C^{\infty}$ foliation, and hence $E^{s}\oplus E^{u}$ is $C^{\infty}$. If $f$ is Anosov, smoothness of $E^{s}\oplus E^{u}$ follows from [45, Theorem 1.3].
@@
To complete the proof of Theorem A, it remains to establish the rigidity result in the accessible case. In this setting $f$ is ergodic, hence topologically transitive. By (3) and (2) in Theorem A, $f$ is center isometric and $E^{s}\oplus E^{u}$ is $C^{\infty}$. Corollary 3.4 then yields that $E^{c}$ is $C^{1}$. The $C^{\infty}$ classification follows from the next proposition.
@@
  • an isometric extension of an Anosov diffeomorphism;

### 16 Apr 2026 v2

@@ 4.1. Adapted charts and templates
    let $F_{x}:=(\imath_{f(x)})^{-1}\circ f\circ\imath_{x}=(F_{x,1},F_{x,2},F_{x,3})$; then, there exist quadratic polynomials $P_{x}^{s}(z)=\alpha_{x}^{\mathrm{FH}}z+\beta_{x}^{s}z^{2}$, $P_{x}^{u}(z)=\alpha_{x}^{\mathrm{FH}}z+\beta_{x}^{u}z^{2}$, such that for $\xi\in(-1,1)$ and $\eta\in(-(\lambda_{x}^{u})^{-1},(\lambda_{x}^{u})^{-1})$, we have
@@
    (4) $\alpha_{x}^{\mathrm{FH}}:=\partial_{13}F_{x,2}(0,0,0),\quad\forall\,x\in M.$
    We emphasize that the notation $\alpha_{x}^{\mathrm{FH}}$ here is unrelated to the contact form $\alpha$ introduced in the previous section.
@@
To obtain the desired polynomials as stated in Proposition 4.2, two additional cohomological equations must be solved. We focus on the case of $r_{x}^{u}$; the other case follows similarly. We decompose $r_{x}^{u}(\xi)$ as $r_{x}^{u}(\xi)=P_{x}^{u}(\xi)+\hat{r}_{x}^{u}(\xi)$, where $P_{x}^{u}(\xi)=\alpha_{x}^{\mathrm{FH}}\xi+\beta_{x}^{u}\xi^{2}$ and $\hat{r}_{x}^{u}(\xi)=O(\xi^{3})$. As in [28, Claim 3.5], the required chart adjustment involves solving the following cohomological equation for $u$:
@@ 4.2. The Foulon–Hasselblatt cocycle
We define the (twisted) Foulon–Hasselblatt cocycle $\alpha^{\mathrm{FH}}$ to be the map
$M\times\mathbb{Z}\ni(x,n)\mapsto\alpha_{x}^{\mathrm{FH}}(n):=\left\{\begin{array}[]{ll}\sum_{\ell=0}^{n-1}\lambda_{x}(\ell)\alpha_{f^{\ell}(x)}^{\mathrm{FH}},&n\geq 0,\\ -\sum_{\ell=1}^{-n}\lambda_{x}(-\ell)\alpha_{f^{-\ell}(x)}^{\mathrm{FH}},&n<0,\end{array}\right$.
with $\alpha_{x}^{\mathrm{FH}}$ as in (4). Note that the twist is given by the cocycle $\lambda$ defined above.
@@
The Foulon–Hasselblatt cocycle $\alpha^{\mathrm{FH}}$ is a twisted coboundary if
$\alpha_{x}^{\mathrm{FH}}=\lambda_{x}\beta(f(x))-\beta(x),\quad\forall\,x\in M,$
@@
The twisted cohomology class of the Foulon–Hasselblatt cocycle is independent of the choice of adapted charts. In other words, if $\{\tilde{\imath}_{x}\}_{x\in M}$ is another family of adapted charts, and $\tilde{\alpha}^{\mathrm{FH}}$ denotes the associated Foulon–Hasselblatt cocycle, then the difference $\tilde{\alpha}^{\mathrm{FH}}-\alpha^{\mathrm{FH}}$ is a twisted coboundary, i.e.,
$\tilde{\alpha}_{x}^{\mathrm{FH}}-\alpha_{x}^{\mathrm{FH}}=\lambda_{x}\beta(f(x))-\beta(x),\quad\forall\,x\in M,$
@@
$\begin{bmatrix}\partial_{1}\tilde{F}_{x,1}&\partial_{2}\tilde{F}_{x,1}\\ \partial_{1}\tilde{F}_{x,2}&\partial_{2}\tilde{F}_{x,2}\end{bmatrix}(0,0,\eta)$ $=\begin{bmatrix}\lambda_{x}^{s}&0\\ \tilde{\alpha}_{x}^{\mathrm{FH}}\eta+\tilde{\beta}_{x}^{s}\eta^{2}&\lambda_{x}^{c}\end{bmatrix}$.
@@
$=(\tilde{\alpha}_{x}^{\mathrm{FH}}+\lambda_{x}^{c}\tau_{x})\xi\eta+H.O.T.$
@@
$\Phi_{f(x)}\circ\tilde{F}_{x}(\xi,0,\eta)$ $=\Phi_{f(x)}(\lambda_{x}^{s}\xi+H.O.T.,\alpha_{x}^{\mathrm{FH}}\xi\eta+H.O.T.,\lambda_{x}^{u}\eta+H.O.T.)$
$=(*,(\alpha_{x}^{\mathrm{FH}}+\tau_{f(x)}\lambda_{x}^{s}\lambda_{x}^{u})\xi\eta+H.O.T.,*).$
@@
(7) $\tilde{\alpha}_{x}^{\mathrm{FH}}-\alpha_{x}^{\mathrm{FH}}=\lambda_{x}^{s}\lambda_{x}^{u}\tau_{f(x)}-\lambda_{x}^{c}\tau_{x}=\lambda_{x}\beta(f(x))-\beta(x),$
@@
$\alpha_{x}^{\mathrm{FH}}=\lambda_{x}\beta(f(x))-\beta(x),\quad\forall\,x\in M.$
We look for a family $\{\tilde{\imath}_{x}\}_{x\in M}$ such that the associated family of (quadratic) polynomials $\tilde{P}_{x}^{*}$ have a vanishing linear coefficient, i.e., $\tilde{\alpha}_{x}^{\mathrm{FH}}=0$. As above, we let $\Phi_{x}:=\tilde{\imath}_{x}^{-1}\circ\imath_{x}\colon(\xi,t,\eta)\mapsto(\tilde{\xi}_{x},\tilde{t}_{x},\tilde{\eta}_{x})$ be the change of coordinates; by (7), we then have
$\tilde{\alpha}_{x}^{\mathrm{FH}}=\alpha_{x}^{\mathrm{FH}}+\lambda_{x}\lambda_{f(x)}^{c}\partial_{13}\tilde{t}_{f(x)}(0,0,0)-\lambda_{x}^{c}\partial_{13}\tilde{t}_{x}(0,0,0).$
To ensure that $\tilde{\alpha}_{x}^{\mathrm{FH}}=0$, it suffices to solve $\lambda_{x}^{c}\partial_{13}\tilde{t}_{x}(0,0,0)=\beta(x)$; we can then take
@@
The non-triviality of the Foulon–Hasselblatt cocycle $\alpha^{\mathrm{FH}}$ is an obstruction to the regularity of the distribution $E^{s}\oplus E^{u}$.
@@
If $E^{s}\oplus E^{u}$ is $C^{1}$ along the leaves of the unstable foliation $\mathcal{W}^{u}$, then $\alpha^{\mathrm{FH}}$ is a twisted coboundary. The same conclusion holds if $E^{s}\oplus E^{u}$ is $C^{1}$ along the leaves of the stable foliation $\mathcal{W}^{s}$.
@@
$\alpha_{x}^{\mathrm{FH}}$ $=\lambda_{x}^{s}\lambda_{x}^{u}(\mathcal{T}_{f(x)}^{s})^{\prime}(0)-\lambda_{x}^{c}(\mathcal{T}_{x}^{s})^{\prime}(0)$
@@ 4.3. Bootstrap of regularity
Note that $f$ is dynamically coherent and center bunched, hence the distribution $E^{s}\oplus E^{u}$ is uniformly $C^{1}$ along $\mathcal{W}^{c}$. It therefore suffices to show that $E^{s}\oplus E^{u}$ is uniformly $C^{1}$ along the local manifolds $\mathcal{W}^{s}_{\mathrm{loc}}(x)$ and $\mathcal{W}^{u}_{\mathrm{loc}}(x)$ for all $x\in M$. By construction, this is equivalent to proving that the templates $\mathcal{T}^{s}_{x}$ and $\mathcal{T}^{u}_{x}$ are uniformly $C^{1}$. We prove this for $\mathcal{T}^{s}_{x}$; the argument for $\mathcal{T}^{u}_{x}$ is analogous.
@@
$\alpha_{x}^{\mathrm{FH}}=\lambda_{x}^{s}\lambda_{x}^{u}(\mathcal{T}_{f(x)}^{s})^{\prime}(0)-\lambda_{x}^{c}(\mathcal{T}_{x}^{s})^{\prime}(0)=\lambda_{x}\beta(f(x))-\beta(x)$
@@
$\alpha^{\prime}_{x}:=\gamma(x)\alpha_{x}^{\mathrm{FH}}=\gamma(x)(\lambda_{x}\beta(f(x))-\beta(x))=\beta(f(x))\gamma(f(x))-\beta(x)\gamma(x)$
@@
By Lemma 4.11, the distribution $E^{s}\oplus E^{u}$ is $C^{1}$. This implies that $x\mapsto\alpha_{x}^{\mathrm{FH}}$ is a twisted coboundary by Lemma 4.10. Then by Lemma 4.9, up to replacing the charts $\{\imath_{x}\}_{x\in M}$ by new charts $\{\tilde{\imath}_{x}\}_{x\in M}$, we can assume that the Foulon–Hasselblatt cocycle vanishes identically. In other words, for all $x\in M$, the degree-one coefficient $\alpha_{x}^{\mathrm{FH}}=(P_{x}^{s})^{\prime}(0)=0$ of the polynomial $P_{x}^{s}$ vanishes; consequently, $P_{x}^{s}(\eta)=\beta_{x}^{s}\eta^{2}$. By (5), for $x\in M$ and $\eta\in(-1,1)$, we have
@@ 5.3. Smoothness of $E^{s}\oplus E^{u}$
We then show that if $E^{s}\oplus E^{u}$ is Lipschitz then it is actually $C^{\infty}$. Recall that we have fixed a non-trivial $f$-invariant open accessibility class $U\subsetneq M$ of $f$. By applying Proposition 4.12, in order to show that $E^{s}\oplus E^{u}$ is $C^{\infty}$ on $U$, it suffices to check that $f_{\bar{U}}$ is $\infty$-bunched.
@@
The center Lyapunov exponent $\chi^{c}(\nu)$ vanishes for every $f$-invariant ergodic measure $\nu$ supported on $\bar{U}$. In particular, the map $f_{\bar{U}}$ is $\infty$-bunched.
@@
Let $x\in\bar{U}$ be a $\nu$-typical point. If $x\in U$, then by Corollary 3.3, the center Lyapunov exponent satisfies
@@ 6.1. Proof of Theorem A
For (3), Proposition 5.1 implies that either $f$ is accessible or $E^{s}\oplus E^{u}$ is integrable. Note that $\lambda^{c}_{x}$ is a coboundary if and only if $f$ is center isometric, i.e., there is a continuous metric such that $\lambda^{c}_{x}=1$, for all $x\in M$. In the accessible case, $\lambda^{c}_{x}$ is a coboundary by Corollary 3.3. If $E^{s}\oplus E^{u}$ is integrable, then by Proposition 2.7 the map $f$ is either center isometric or topologically conjugate to an Anosov automorphism of ${\mathbb{T}}^{3}$. In the latter case $f$ is Anosov, and $\lambda^{c}$ is cohomologous to a constant by [32, Theorem 1.1].
For (1), center isometric partially hyperbolic systems are dynamically coherent by [57, Theorem 7.5], while Anosov diffeomorphisms on ${\mathbb{T}}^{3}$ are always dynamically coherent [55].
For (2), if $f$ is center isometric and accessible, then $E^{s}\oplus E^{u}$ is $C^{\infty}$ by Proposition 4.12. If $E^{s}\oplus E^{u}$ integrates to a foliation $\mathcal{W}^{su}$, then the leaves of $\mathcal{W}^{su}$ are $C^{\infty}$ by Lemma 2.10. Since $f$ is $\infty$-bunched, the holonomy of $\mathcal{W}^{su}$ along $\mathcal{W}^{c}$ is uniformly $C^{\infty}$. Journé’s lemma [44] then yields that $\mathcal{W}^{su}$ is a $C^{\infty}$ foliation, and hence $E^{s}\oplus E^{u}$ is $C^{\infty}$. If $f$ is Anosov, smoothness of $E^{s}\oplus E^{u}$ follows from [47, Theorem 1.3].
@@
To complete the proof of Theorem A, it remains to establish the rigidity result in the accessible case. In this setting $f$ is ergodic, hence topologically transitive. By (3) and (2) in Theorem A, $f$ is center isometric and $E^{s}\oplus E^{u}$ is $C^{\infty}$. Corollary 3.5 then yields that $E^{c}$ is $C^{\infty}$ and $f$ preserves a $C^{\infty}$ contact form. The $C^{\infty}$ classification follows from the next proposition.
@@
  • an isometric extension of a volume-preserving Anosov diffeomorphism on ${\mathbb{T}}^{2}$;
@@
Finally, if $f$ is $C^{\infty}$-conjugate to an isometric extension of an Anosov diffeomorphism, then $M$ is either a torus or a Heisenberg manifold. We prove by contradiction that $M\neq{\mathbb{T}}^{3}$. Suppose instead that $M={\mathbb{T}}^{3}$, then $\mathcal{W}^{c}$ is a trivial bundle over ${\mathbb{T}}^{2}$. Let $S\subset{\mathbb{T}}^{3}$ be a global section transverse to $\mathcal{W}^{c}$, and let $\beta$ be the $C^{\infty}$ contact form preserved by $f$. Arguing as in Section 5.4, Lemma 2.4 implies that $d\beta|_{TS}$ is nondegenerate. Consequently,
$\int_{S}d\beta\neq 0,$
which contradicts Stokes’ theorem.